OpenStudy (anonymous):

x2 + y2 = 13 where y = 2 x = _

OpenStudy (anonymous):

x2+2*2=13 x2+4=13 Can you take it from here?

OpenStudy (anonymous):

@pricelesslove247

OpenStudy (anonymous):

Im sorry I still dont get it

OpenStudy (anonymous):

okay... what are you misunderstanding? Do you not understand how I got that?

OpenStudy (anonymous):

Yeah i dont understand how you got that

OpenStudy (anonymous):

Y=2, yes? @pricelesslove247

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay, so plug it into your equation. Wherever it says "Y", put a 2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so what do you have

OpenStudy (anonymous):

x2 + 2 X 2 = 13

OpenStudy (anonymous):

You cause use " * " for times

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

solve it. what is 2*2? put your answer in that slot but keep the rest the same

OpenStudy (anonymous):

3

OpenStudy (anonymous):

How old are you?

OpenStudy (anonymous):

that answer is 4 but is the final answer 3

OpenStudy (anonymous):

9 divided by 2 is 4.5 If you came up with the final answer on your own, why did you post it here?